\begin{thebibliography}{25}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Abdelali et~al.(2016)Abdelali, Darwish, Durrani, and
  Mubarak}]{DBLP:conf/naacl/AbdelaliDDM16}
Ahmed Abdelali, Kareem Darwish, Nadir Durrani, and Hamdy Mubarak. 2016.
\newblock Farasa: {A} fast and furious segmenter for arabic.
\newblock In \emph{Proceedings of the Demonstrations Session, {NAACL} {HLT}
  2016, The 2016 Conference of the North American Chapter of the Association
  for Computational Linguistics: Human Language Technologies, San Diego
  California, USA, June 12-17, 2016}, pages 11--16.

\bibitem[{Almahairi et~al.(2016)Almahairi, Cho, Habash, and
  Courville}]{AlmahairiCHC16}
Amjad Almahairi, Kyunghyun Cho, Nizar Habash, and Aaron~C. Courville. 2016.
\newblock First result on {A}rabic neural machine translation.
\newblock \emph{Computing Research Repository}, arXiv:1606.02680.
\newblock Version 1.

\bibitem[{Bahdanau et~al.(2015)Bahdanau, Cho, and Bengio}]{BahdanauCB14}
Dzmitry Bahdanau, Kyunghyun Cho, and Yoshua Bengio. 2015.
\newblock Neural machine translation by jointly learning to align and
  translate.
\newblock In \emph{Proceedings of the International Conference on Learning
  Representations (ICLR)}.

\bibitem[{Bojar et~al.(2017)Bojar, Chatterjee, Federmann, Graham, Haddow,
  Huang, Huck, Koehn, Liu, Logacheva, Monz, Negri, Post, Rubino, Specia, and
  Turchi}]{bojar-EtAl:2017:WMT1}
Ond\v{r}ej Bojar, Rajen Chatterjee, Christian Federmann, Yvette Graham, Barry
  Haddow, Shujian Huang, Matthias Huck, Philipp Koehn, Qun Liu, Varvara
  Logacheva, Christof Monz, Matteo Negri, Matt Post, Raphael Rubino, Lucia
  Specia, and Marco Turchi. 2017.
\newblock Findings of the 2017 conference on machine translation (wmt17).
\newblock In \emph{Proceedings of the Second Conference on Machine Translation,
  Volume 2: Shared Task Papers}, pages 169--214, Copenhagen, Denmark.
  Association for Computational Linguistics.

\bibitem[{Cho et~al.(2014)Cho, van Merrienboer, Gulcehre, Bahdanau, Bougares,
  Schwenk, and Bengio}]{cho-EtAl:2014:EMNLP2014}
Kyunghyun Cho, Bart van Merrienboer, Caglar Gulcehre, Dzmitry Bahdanau, Fethi
  Bougares, Holger Schwenk, and Yoshua Bengio. 2014.
\newblock Learning phrase representations using rnn encoder--decoder for
  statistical machine translation.
\newblock In \emph{Proceedings of the 2014 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 1724--1734, Doha, Qatar.
  Association for Computational Linguistics.

\bibitem[{Devlin et~al.(2014)Devlin, Zbib, Huang, Lamar, Schwartz, and
  Makhoul}]{devlin-EtAl}
Jacob Devlin, Rabih Zbib, Zhongqiang Huang, Thomas Lamar, Richard Schwartz, and
  John Makhoul. 2014.
\newblock Fast and robust neural network joint models for statistical machine
  translation.
\newblock In \emph{Proceedings of the 52nd Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1370--1380,
  Baltimore, Maryland. Association for Computational Linguistics.

\bibitem[{Dyer et~al.(2013)Dyer, Chahuneau, and
  Smith}]{dyer-chahuneau-smith:2013:NAACL-HLT}
Chris Dyer, Victor Chahuneau, and Noah~A. Smith. 2013.
\newblock A simple, fast, and effective reparameterization of ibm model 2.
\newblock In \emph{Proceedings of the 2013 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 644--648, Atlanta, Georgia. Association for
  Computational Linguistics.

\bibitem[{Heafield et~al.(2013)Heafield, Pouzyrevsky, Clark, and
  Koehn}]{heafield-EtAl:2013:Short}
Kenneth Heafield, Ivan Pouzyrevsky, Jonathan~H. Clark, and Philipp Koehn. 2013.
\newblock Scalable modified kneser-ney language model estimation.
\newblock In \emph{Proceedings of the 51st Annual Meeting of the Association
  for Computational Linguistics (Volume 2: Short Papers)}, pages 690--696,
  Sofia, Bulgaria. Association for Computational Linguistics.

\bibitem[{Junczys{-}Dowmunt et~al.(2016)Junczys{-}Dowmunt, Dwojak, and
  Hoang}]{Junczys-Dowmunt16c}
Marcin Junczys{-}Dowmunt, Tomasz Dwojak, and Hieu Hoang. 2016.
\newblock Is neural machine translation ready for deployment? {A} case study on
  30 translation directions.
\newblock In \emph{Proceedings of the 13th International Workshop on Spoken
  Language Translation}.

\bibitem[{Junczys-Dowmunt et~al.(2018)Junczys-Dowmunt, Grundkiewicz, Dwojak,
  Hoang, Heafield, Neckermann, Seide, Germann, Aji, Bogoychev, Martins, and
  Birch}]{junczys2018marian}
Marcin Junczys-Dowmunt, Roman Grundkiewicz, Tomasz Dwojak, Hieu Hoang, Kenneth
  Heafield, Tom Neckermann, Frank Seide, Ulrich Germann, Alham~Fikri Aji,
  Nikolay Bogoychev, André F.~T. Martins, and Alexandra Birch. 2018.
\newblock Marian: Fast neural machine translation in c++.
\newblock \emph{Computing Research Repository}, arXiv:1804.00344.
\newblock Version 3.

\bibitem[{Kalchbrenner and Blunsom(2013)}]{kalchbrenner-blunsom:2013:EMNLP}
Nal Kalchbrenner and Phil Blunsom. 2013.
\newblock Recurrent continuous translation models.
\newblock In \emph{Proceedings of the 2013 Conference on Empirical Methods in
  Natural Language Processing}, pages 1700--1709, Seattle, Washington, USA.
  Association for Computational Linguistics.

\bibitem[{Kneser and Ney(1995)}]{Kneser}
Reinhard Kneser and Hermann Ney. 1995.
\newblock Improved backing-off for m-gram language modeling.
\newblock In \emph{Proceedings of the IEEE International Conference on
  Accoustics, Speech and Signal Processing}, volume~1.

\bibitem[{Koehn et~al.(2007)Koehn, Hoang, Birch, Callison-Burch, Federico,
  Bertoldi, Cowan, Shen, Moran, Zens, Dyer, Bojar, Constantin, and
  Herbst}]{Koehn2007}
Philipp Koehn, Hieu Hoang, Alexandra Birch, Chris Callison-Burch, Marcello
  Federico, Nicola Bertoldi, Brooke Cowan, Wade Shen, Christine Moran, Richard
  Zens, Christopher~J. Dyer, Ond\v{r}ej Bojar, Alexandra Constantin, and Evan
  Herbst. 2007.
\newblock Moses: Open source toolkit for statistical machine translation.
\newblock In \emph{Proceedings of the 45th Annual Meeting of the Association
  for Computational Linguistics Companion Volume Proceedings of the Demo and
  Poster Sessions}, pages 177--180, Prague, Czech Republic. Association for
  Computational Linguistics.

\bibitem[{Koehn and Knowles(2017)}]{koehn-knowles:2017:NMT}
Philipp Koehn and Rebecca Knowles. 2017.
\newblock Six challenges for neural machine translation.
\newblock In \emph{Proceedings of the First Workshop on Neural Machine
  Translation}, pages 28--39, Vancouver. Association for Computational
  Linguistics.

\bibitem[{Och(2003)}]{Och2003b}
Franz~Josef Och. 2003.
\newblock Minimum error rate training in statistical machine translation.
\newblock In \emph{Proceedings of the 41st Annual Meeting of the Association
  for Computational Linguistics}, pages 160--167.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and Zhu}]{Papineni}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock Bleu: a method for automatic evaluation of machine translation.
\newblock In \emph{Proceedings of 40th Annual Meeting of the Association for
  Computational Linguistics}, pages 311--318, Philadelphia, Pennsylvania, USA.
  Association for Computational Linguistics.

\bibitem[{Rafalovitch and Dale(2009)}]{Rafalovitch}
Alexandre Rafalovitch and Robert Dale. 2009.
\newblock United nations general assembly resolutions: A six-language parallel
  corpus.
\newblock In \emph{Proceedings of the MT Summit XII}, pages 292--299.

\bibitem[{Sennrich et~al.(2017)Sennrich, Firat, Cho, Birch, Haddow, Hitschler,
  Junczys-Dowmunt, L\"{a}ubli, Miceli~Barone, Mokry, and
  Nadejde}]{sennrich-EtAl:2017:EACLDemo}
Rico Sennrich, Orhan Firat, Kyunghyun Cho, Alexandra Birch, Barry Haddow,
  Julian Hitschler, Marcin Junczys-Dowmunt, Samuel L\"{a}ubli, Antonio~Valerio
  Miceli~Barone, Jozef Mokry, and Maria Nadejde. 2017.
\newblock Nematus: a toolkit for neural machine translation.
\newblock In \emph{Proceedings of the Software Demonstrations of the 15th
  Conference of the European Chapter of the Association for Computational
  Linguistics}, pages 65--68, Valencia, Spain. Association for Computational
  Linguistics.

\bibitem[{Sennrich et~al.(2016{\natexlab{a}})Sennrich, Haddow, and
  Birch}]{sennrich-haddow-birch:2016:P16-11}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016{\natexlab{a}}.
\newblock Improving neural machine translation models with monolingual data.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 86--96, Berlin,
  Germany. Association for Computational Linguistics.

\bibitem[{Sennrich et~al.(2016{\natexlab{b}})Sennrich, Haddow, and
  Birch}]{sennrich-haddow-birch:2016:P16-12}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016{\natexlab{b}}.
\newblock Neural machine translation of rare words with subword units.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1715--1725,
  Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Sutskever et~al.(2014)Sutskever, Vinyals, and Le}]{NIPS2014_5346}
Ilya Sutskever, Oriol Vinyals, and Quoc~V Le. 2014.
\newblock Sequence to sequence learning with neural networks.
\newblock In Z.~Ghahramani, M.~Welling, C.~Cortes, N.~D. Lawrence, and K.~Q.
  Weinberger, editors, \emph{Advances in Neural Information Processing Systems
  27}, pages 3104--3112. Curran Associates, Inc.

\bibitem[{Tiedemann(2012)}]{TIEDEMANN12.463}
Jörg Tiedemann. 2012.
\newblock Parallel data, tools and interfaces in opus.
\newblock In \emph{Proceedings of the Eight International Conference on
  Language Resources and Evaluation (LREC'12)}, Istanbul, Turkey. European
  Language Resources Association (ELRA).

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{NIPS2017_7181}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, \L~ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock Attention is all you need.
\newblock In I.~Guyon, U.~V. Luxburg, S.~Bengio, H.~Wallach, R.~Fergus,
  S.~Vishwanathan, and R.~Garnett, editors, \emph{Advances in Neural
  Information Processing Systems 30}, pages 5998--6008. Curran Associates, Inc.

\bibitem[{Wołk and Marasek(2014)}]{WOLK2014126}
Krzysztof Wołk and Krzysztof Marasek. 2014.
\newblock Building subject-aligned comparable corpora and mining it for truly
  parallel sentence pairs.
\newblock \emph{Procedia Technology}, 18:126 -- 132.
\newblock International workshop on Innovations in Information and
  Communication Science and Technology, IICST 2014, 3-5 September 2014, Warsaw,
  Poland.

\bibitem[{Ziemski et~al.(2016)Ziemski, Junczys-Dowmunt, and
  Pouliquen}]{ZIEMSKI16.1195}
Michał Ziemski, Marcin Junczys-Dowmunt, and Bruno Pouliquen. 2016.
\newblock The united nations parallel corpus v1.0.
\newblock In \emph{Proceedings of the Tenth International Conference on
  Language Resources and Evaluation (LREC 2016)}, Paris, France. European
  Language Resources Association (ELRA).

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{emnlp2018}
\usepackage{times}
\usepackage{latexsym}
\usepackage{amsmath}
\usepackage{graphicx}
\usepackage{multirow}
\usepackage{url}

\aclfinalcopy 

\title{A Recipe for Arabic-English Neural Machine Translation}

\author{Abdullah Alrajeh \\
  National Center for AI and Big Data Technology \\
  King Abdulaziz City for Science and Technology (KACST) \\
  Riyadh, Saudi Arabia \\
  {\tt asrajeh@kacst.edu.sa}}

\date{}

\begin{document}
\maketitle
\begin{abstract}
In this paper, we present a recipe for building a good Arabic-English
neural machine translation. We compare neural systems with traditional 
phrase-based systems using various parallel corpora including UN, ISI 
and Ummah. We also investigate the importance of special preprocessing 
of the Arabic script. The presented results are based on test sets from 
NIST MT 2005 and 2012. The best neural system produces a gain of +13 BLEU 
points compared to an equivalent simple phrase-based system in NIST MT12
test set. Unexpectedly, we find that tuning a model trained on the whole 
data using a small high quality corpus like Ummah gives a substantial 
improvement (+3 BLEU points). We also find that training a neural system with
a small Arabic-English corpus is competitive to a traditional phrase-based
system.

\end{abstract}

\section{Introduction}
Neural networks succeed to show impressive results as part of a statistical
machine translation (SMT) system in the work of \citet{devlin-EtAl}. Since then, 
the research shifted more towards an end-to-end approach. Currently, neural machine 
translation (NMT) has become the dominant approach in the field achieving 
state-of-the-art results in many translation tasks \cite{bojar-EtAl:2017:WMT1}.

\citet{Junczys-Dowmunt16c} investigate 30 translation directions using the UN corpus 
(around 335M words). The experiments, based on test sets from the same corpus, show 
that NMT is superior to the traditional approach (i.e. phrase-based SMT). One of the 
investigated task is translation between Arabic and English but without special 
preprocessing for Arabic. Large improvement is observed (around 3 BLEU points) over 
phrase-based SMT in both directions.

\citet{AlmahairiCHC16} compared a neural system against phrase-based one in Arabic-English
translation task and found them to be comparable based on NIST 2005 test set. It is also observed 
that NMT is superior to SMT in an out-of-domain test set. In all cases, preprocessing
of Arabic script did improve the translation quality.

In this paper, we further investigate Arabic to English translation using several corpora
including Ummah, ISI, UN and many others. We compare the performance of NMT against 
phrase-based SMT. In our experiments, we applied Arabic preprocessing, which includes 
normalization and tokenization, to see its impact on both NMT and SMT systems. Our results 
is based on NIST MT sets for the year 2005, 2006 and 2012.

In the next section, we give a brief introduction to neural machine translation. Section 3
lists the parallel corpora will be used and show some statistics. Section 4 presents our SMT 
and NMT experiments followed by the conclusion.

\section{Neural Machine Translation}
\citet{kalchbrenner-blunsom:2013:EMNLP} set the foundation of neural machine translation by
proposing an end-to-end encoder-decoder approach. Convolutional neural networks (CNN) are used
to encode a source sentence then generates its translation by recurrent neural networks (RNN).

Long sentences propose a challenge for RNN where there are long distance reordering.
\citet{NIPS2014_5346} develop sequence-to-sequence models that use RNN for both encoding 
and decoding. Standard RNN units are replaced with Long Short-Term Memory (LSTM) units to 
capture long-term dependencies. \citet{cho-EtAl:2014:EMNLP2014} introduce Gated Recurrent 
Unit (GRU) which is simpler than LSTM. 

In the previous work, a source sentence is encoded into a fixed-length vector which is
a bottleneck holding NMT from being competitive to SMT particularly in long sentences.
\citet{BahdanauCB14} introduce the powerful attention mechanism that allows the decoder
to focus on different words while translating.

These advancements and other such as byte pair encoding \cite{sennrich-haddow-birch:2016:P16-12},
to achieve open vocabulary NMT, pave the way for new state-of-the-art translation systems.

Mathematically, the probability of a translation sentence (${\bf y}:y_1,\dots,y_I$) of 
an input sentence (${\bf x}:x_1,\dots,x_J$) is computed as follows:
\begin{align}
\centering
&p({\bf y}|{\bf x}) = \prod^I_{i=1} p(y_i|y_1,\dots,y_{i-1},{\bf x}) \\
&p(y_i|y_1,\dots,y_{i-1},{\bf x}) = g(y_{i-1},s_i,c_i)
\end{align}

where $s_i=f(s_{i-1},y_{i-1},c_i)$ is a hidden state in the RNN decoder while $c_i$ is the context
vector computed from all hidden states ($h$) in the RNN encoder as follows: 
\begin{align}
\centering
&c_i=\sum^J_{j=1} \alpha_{ij} h_j \\
&\alpha_{ij} = \frac{\exp(a(s_{i-1},h_j))}{\sum^J_{k=1}\exp(a(s_{i-1},h_k))}
\end{align}

where $h_j=f(h_{j-1},x_j)$, $\alpha_{ij}$ is its weight and $a(.)$ is an alignment model shows 
the importance of the input word $j$ in translating the output word $i$. This mechanism allows the 
decoder to just pay attention to the related input words. Note that, the function $f$ that produces 
the next hidden state in the encoder and decoder can be defined as LSTM or GRU.

Usually, an input sentence is encoded by a forward RNN but a backward RNN, that reads the sentence in a 
reverse order, is found to improve the performance \cite{NIPS2014_5346}. A bidirectional RNN also has
been successful \cite{BahdanauCB14}. It reads the sentence in both directions then concatenates the 
forward and backword hidden states as follows:
\begin{equation}
\centering
h_j=[\overrightarrow{h}^T_j;\overleftarrow{h}^T_j]^T
\end{equation}

\section{Corpora}
There are many parallel corpora available for building Arabic-English translation systems. The UN 
corpus\footnote{https://conferences.unite.un.org/uncorpus} is an obvious choice for many researchers 
and will be used in our experiments. It is composed of parliamentary documents of the United Nations 
since 1990 for Arabic, Chinese, English, French, Russian, and Spanish \cite{ZIEMSKI16.1195}.

There are also 11 LDC\footnote{http://ldc.upenn.edu/} corpora have been selected. These include Ummah and ISI
with catalogue numbers LDC2004T18 and LDC2007T08, respectively. Ummah corpus contains news stories while ISI
was extracted automatically from Arabic Gigaword and English Gigaword. The rest are mostly from GALE project
with catalogue numbers LDC2004T17, LDC2005T05, LDC2008T09, LDC2009T09, LDC2013T10, LDC2013T14, LDC2015T05, 
LDC2015T07 and LDC2015T19.

Besides that, we used all Arabic-English corpora available on OPUS\footnote{http://opus.nlpl.eu/} website
\cite{TIEDEMANN12.463,Rafalovitch,WOLK2014126}. We exclude MultiUN because we already have larger version. 
OpenSubtitles and Tanzil are also excluded due to their low quality.

Table~\ref{ar-en-corpora} shows statistics of all corpora. The total number of all English words is close to 
half a billion words.

\begin{table}[h]
\centering
\resizebox{\columnwidth}{!}{%
\begin{tabular}{|c|c|c|cc|} \hline
No. & Corpus & Sentences & \multicolumn{2}{c|}{Ar-En words} \\ \hline
1   & Ummah		 & 80k	 & 2.3m  & 2.9m  \\
2   & ISI		 & 1.1m  & 28.9m & 30.8m \\
3   & \small{LDC2004T17} & 19k & 441k & 581k \\
4   & \small{LDC2005T05} & 5k & 106k & 135k \\
5   & \small{LDC2008T09} & 3k & 55k & 68k \\       
6   & \small{LDC2009T09} & 10k & 145k & 198k \\       
7   & \small{LDC2013T10} & 8k & 182k & 240k \\
8   & \small{LDC2013T14} & 5k & 89k & 124k \\              
9   & \small{LDC2015T05} & 18k & 285k & 379k \\       
10  & \small{LDC2015T07} & 20k & 330k & 440k \\       
11  & \small{LDC2015T19} & 6K & 156k & 210k \\
12  & OPUS 		 & 639k & 13.8m & 13.8m \\       
13  & UN 		 & 185m & 398m & 448m \\ \hline
\end{tabular}
}
\caption{Statistics of all Arabic-English corpora (m: million, k: thousand).}
\label{ar-en-corpora}
\end{table}

\section{Experiments}
We present SMT and NMT results on Arabic-English based on NIST MT sets for 
the year 2005, 2006 and 2012 (see Table~\ref{nist-sets}). As commonly used 
in machine translation, we evaluated the translation performance by BLEU 
score \citep{Papineni}.

\begin{table}[h]
\centering
\resizebox{\columnwidth}{!}{%
\begin{tabular}{|c|cc|cc|} \hline
          & \multicolumn{2}{c|}{Ar-En Sentences} & \multicolumn{2}{c|}{Ar-En words} \\ \hline
MT06 (dev)  & 1797   & 1797 & 42k & 54k  \\
MT05 (test) & 1056   & 4224 & 26k & 130k \\
MT12 (test) & 1378   & 5512 & 35k & 191k \\ \hline
\end{tabular}
}
\caption{Statistics of NIST MT sets.}
\label{nist-sets}
\end{table}

The systems are 
trained on different datasets ranging from small to very large. Training corpora
in Table~\ref{ar-en-corpora} are grouped into 4 sets:
\begin{itemize}
\item Set A: Ummah corpus
\item Set B: Ummah, ISI and LDC2004T17
\item Set C: all corpora except UN
\item Set D: all corpora
\end{itemize}

The reasons for this setting are the following.
Low-resource MT task is a known challenge for NMT \cite{koehn-knowles:2017:NMT}.
We would like to see if this is the case for Arabic-English task (Set A). 
\citet{AlmahairiCHC16} report the first result on Arabic NMT therefore Set B are 
chosen for comparison. Finally, UN corpus might add no benefit \cite{devlin-EtAl}
since it is not the news domain (Set C and D).

{\bf Preprocessing} \, In our experiments, we applied Arabic preprocessing, which includes
normalization and tokenization (ATB scheme), to see its impact on both SMT and NMT systems.
We used Farasa \cite{DBLP:conf/naacl/AbdelaliDDM16}, a fast Arabic segmenter.
The maximum sentence length is 100.

{\bf Phrase-based MT} \, We use Moses toolkit \citep{Koehn2007} with its default settings. 
The language model is a 5-gram built from the English side with interpolation and Kneser-Ney 
smoothing \citep{Kneser} built by KenLM \cite{heafield-EtAl:2013:Short}. Word alignments are 
extracted by {\tt fast\_align} \cite{dyer-chahuneau-smith:2013:NAACL-HLT}. We tune the system using 
MERT technique \cite{Och2003b}. The chosen option for the reordering model is {\tt msd-bidirectional-fe}.

{\bf Neural MT} \, We use Marian, an efficien and fast NMT system written in C++ \cite{junczys2018marian}.
The system has implemented several models. The {\tt s2s} option is chosen which is equivalent to Nematus 
models \cite{sennrich-EtAl:2017:EACLDemo} that are RNN encoder-decoder based with attention mechanism.
The basic training script provided is used. To achieve open vocabulary, we apply byte pair encoding (BPE)
\cite{sennrich-haddow-birch:2016:P16-12} setting the maximum size of the joint Arabic-English vocabulary to 90,000.

Table~\ref{smt} reports BLEU scores of many phrase-based SMT systems trained on various datasets.
Clearly, preprocessing of the Arabic side is important. Substantial gain is observed when the training
data is small (Set A). Note that adding UN corpus to the training data improves the BLEU score as in set D.

\begin{table}[h]
\centering
\resizebox{\columnwidth}{!}{%
\begin{tabular}{|c|l|c|c|c|} \hline
Set & System & MT05 & MT12 & avg\\ \hline
\multirow{2}{*}{A} & baseline 		& 39.49 & 22.50 & 31.00\\
		   & + ar preprocessing  & 42.17 & 31.87 & 37.02\\ \hline
\multirow{2}{*}{B} & baseline  		& 49.71 & 34.25 & 41.98\\ 
                   & + ar preprocessing  & 51.65 & 37.06 & 44.35\\ \hline
\multirow{2}{*}{C} & baseline           & 51.32 & 38.12 & 44.72\\
                   & + ar preprocessing  & 52.76 & 40.80 & 46.78\\ \hline
\multirow{2}{*}{D} & baseline           & 52.57 & 40.02 & 46.30\\
                   & + ar preprocessing  & 53.45 & 41.11 & 47.28\\ \hline
\end{tabular}
}
\caption{BLEU scores of Arabic-English SMT.}
\label{smt}
\end{table}

Table~\ref{nmt} presents NMT systems's performance in BLEU. Compared to Table~\ref{smt}, NMT is
superior to SMT in all cases. The best NMT system produces a gain of +13 BLEU points in NIST MT12 
test set. Unexpectedly, NMT is similar or better than SMT even with a small dataset
(Set A is less than 3 million English words). Note that the gap in BLEU between NMT and SMT increases
with more training data. Arabic preprocessing improves the performance as in Table~\ref{smt} which
indicates that BPE is not sufficient. We find that tuning a model trained on the whole data using
a high quality corpus like Ummah (Set A) gives us a substantial improvement. Finally, 
an independent ensemble of 5 best models boosts the score with +1.5 BLEU.

\begin{table}[h]
\centering
\resizebox{\columnwidth}{!}{%
\begin{tabular}{|c|l|c|c|c|} \hline
Set & System & MT05 & MT12 & avg\\ \hline
\multirow{2}{*}{A} & baseline           & 41.62 & 19.47 & 30.55\\
                   & + ar preprocessing & 44.15 & 31.86 & 38.01\\ \hline
\multirow{2}{*}{B} & baseline           & 53.27 & 38.19 & 45.73\\
                   & + ar preprocessing & 54.69 & 40.07 & 47.38\\ \hline
\multirow{2}{*}{C} & baseline           & 57.02 & 45.73 & 51.38\\
                   & + ar preprocessing & 58.35 & 47.04 & 52.70\\ \hline
\multirow{2}{*}{D} & baseline           & 57.31 & 45.81 & 51.56\\
                   & + ar preprocessing & 58.43 & 47.74 & 53.09\\
                   & + tuning		& 61.26 & 52.53 & 56.90\\ 
                   & + ensemble of 5 	& 62.98 & 54.27 & 58.63\\ \hline
\end{tabular}
}
\caption{BLEU scores of Arabic-English NMT.}
\label{nmt}
\end{table}

Training the whole data on a single GPU took 4 days\footnote{NVidia GTX 1080 Ti, CPU 4.20GHz and Hard disk SSD}.
The disk size of the best model is just 645 MB. It is very compact compared to the phrase table alone in 
SMT which is 8.5 GB.

During the experiments, other models have been tried like {\tt transformer} \cite{NIPS2017_7181}
but no improvement is gained. It is also the case for the joint vocabulary's size.

\section{Conclusion}
We present Arabic to English machine translation using various training datasets.
We compare neural systems with traditional ones (i.e. phrase-based SMT). We also
investigate the importance of special preprocessing of the Arabic script. The systems
are tested on NIST MT 2005 and 2012.

After the experiments, we draw the following conclusions. In both NMT and SMT systems, Arabic preprocessing 
improves the translation quality as found by \citet{AlmahairiCHC16}. Although UN corpus is not in the news 
domain, a gain is observed in both systems. Neural MT is superior to phrase-based MT in all cases.
NMT able to perform very well given a small corpus. Finally, tuning a model trained on the whole data using
a small high quality corpus (i.e. Ummah) gives a substantial improvement. The best NMT system produces a gain 
of +13 BLEU points in NIST MT12 test set.

There are techniques we have not considered in this work but might improve the translation quality such as
back translation \cite{sennrich-haddow-birch:2016:P16-11}.

\bibliography{emnlp2018}
\bibliographystyle{acl_natbib_nourl}

\end{document}